OpenStudy (zenmo):

For what values of r does the function y = e^{rx} satisfy the differential equation 7y'' + 48y' − 7y = 0?

OpenStudy (zenmo):

\[y = e^{rx}, y' = re^{rx}, y'' = r^2e^{rx}\]\[7(r^2e^{rx})+48(re^{rx})-7(e^{rx})=0\]\[7e^{rx}(r^2+7r-1) = 0\]Am I doing it correctly so far?

ganeshie8 (ganeshie8):

Is the middle coefficient 48 or 49 ?

ganeshie8 (ganeshie8):

Other than that you seem to be on right track. Notice that e^x is never 0. So the other factor must be 0 for the overall product to be 0.

OpenStudy (zenmo):

opps, the middle coefficient is 48, but I got the problem now. :)